OpenStudy (anonymous):

pre-cal please help, pretty simple but just can't think. 1. Solve: 3(x-5) ≤x-9

OpenStudy (anonymous):

i think i got it

OpenStudy (anonymous):

what is x though

OpenStudy (anonymous):

It doesn't say but I know this is what you have to do in one step 

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x = 17?

OpenStudy (anonymous):

thats right^^^^

OpenStudy (anonymous):

did you get it

OpenStudy (anonymous):

No, I know x doesn't equal 17

OpenStudy (anonymous):

i suggest use mathway that is all im gonna do

OpenStudy (anonymous):

I did but it doesn't show you the work

OpenStudy (anonymous):

i cant ether srry tho

OpenStudy (welshfella):

3x - 15 <= x - 9 subtract x from both sides 2x - 15 <= -9 now add 15 to both sides then divide both sides by 2

OpenStudy (anonymous):

Ah okay, got it. Thanks

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

Wait, where did you get 15 @welshfella

OpenStudy (welshfella):

oh right i distributed 3 over x - 5 3(x - 5) = 3*x + 3*-5) = 3x - 15

OpenStudy (welshfella):